### v1

@@ Abstract
We investigate the novel electron beam resist Medusa 84 SiH by Allresist GmbH (Germany) for nanostructuring of single crystal diamond and its effects on the spin properties of nitrogen vacancy (NV) centers in nanopillars as prototypes for photonic structures. We find contrast curves comparable to those of resists previously used for this task (Hydrogensilsequioxane FOx). We present a minimum selectivity for diamond etching of 61. Using an adhesion-promoting silicon interlayer enables fabrication yields of up to 98%. We measure times of up to 25 s before and after processing, demonstrating that the manufactured structures are usable for diamond-based quantum sensing.
@@
In addition, diamond is chemically inert, making it a suitable sensor material for harsh chemical environments. However, considering nanofabrication, chemical inertness poses a challenge; obtaining nanostructures typically requires plasma etching of diamond in a high-density plasma in an inductively coupled plasma reactive ion etcher (ICP-RIE). These plasmas contain oxygen and use a high bias to obtain steep sidewalls of diamond nanostructures. Such plasma conditions, however, pose high demands on the mask material used for etching; strong mask erosion might occur and consequently limit the etch depth and aspect ratio of diamond nanostructures. While, metal masks fabricated in lift-off processes can withstand the plasma etching condition, they might lead to non-optimal side walls of the nanostructures. For many previous works, a mask which is obtained directly via electron beam lithography has been used: Hydrogen Silsesquioxane (HSQ)-based electron beam resists are able to withstand harsh conditions, as the chemical composition of the exposed and developed resist closely resembles SiO2 (”spin-on-glass”).
@@
A main product used in literature [4, 5] as well as in our previous work [6, 20, 21, 7] that satisfies these specifications is Flowable Oxide (FOx, DuPont, formerly DowDuPont). In 2023, DuPont notified distributors that production of FOx-related products would cease at the end of 2025. To make nanofabrication routines for diamond quantum technologies future-proof, also in the light of cost-effectiveness (upper four digit $ range per liter with minimum order quantity of one liter and shelf life of 6 months for FOx), we here present a process for a novel HSQ based resist called Medusa 84 SiH (Allresist GmbH), which has recently become commercially available. The active polymer HSQ is manufactured solely by Allresist and is not related to currently commercially available alternatives. In contrast to DuPont’s FOx, a key difference lies in the chosen solvent which is butyl acetate instead of methyl isobutyl ketone (MIBK), where the latter is considered to be possibly carcinogenic to humans, meaning Medusa 84 SiH also comes with the advantage of higher personnel safety during its application. In this work, we use this novel resist to pattern diamond nanopillars as they are typical for scanning probe sensing using color centers as a prototypical nanostructuring process. The color centers embedded in our nanostructures are individual nitrogen vacancy (NV) centers, which are the most common color centers for quantum sensing.
@@ 2 Nanofabrication Process
To make the following description of our nanofabrication process more readable, we summarize the parameters for devices, materials, and processes in tables in the supplementary material (Tables 1 and 2). In this study, we use three types of samples to evaluate the performance of Medusa 84 SiH. First, we use single crystal, single side polished silicon substrates (Microchemicals), to create accurate contrast curves and examine the writing dose-dependent etch resistivity of Medusa 84 SiH. Second, we use (100)-oriented, single crystal, chemical vapor deposited diamond (Element Six, 3.2 mm 3.2 mm (0.250 mm 0.050 mm)). These diamonds are specified with a moderate surface quality (Ra 30 nm) and less than 1 ppm substitutional nitrogen [N]s. We use this material to develop a consistent nanofabrication process and to study mask erosion. Third, we use a high purity (100)-oriented single crystal electronic grade diamond (Element Six), with dimensions of 4.0 mm 2.0 mm (0.047 mm 0.002 mm) ([N]s 5 ppb, [B] 1 ppb). This diamond has been polished by Delaware Diamond Knives to a surface roughness of Rq = 0.6 nm 0.5 nm (determined via 5 5 m 5 m atomic force microscopy scans (Bruker FastScan Bio)). With this sample, we investigate how Medusa 84 SiH influences NV centers in nanopillars, which are often used in sensing applications.
@@
We here follow published approaches to create shallow NV centers [22, 6, 20]. To obtain coherent NV centers in low-strain material, we remove polishing damage prior to nitrogen implantation by conducting a ”Stress Relief Etch” using an ICP-RIE (Oxford Instruments Plasmalab 100, for etch recipes refer to Table 2) [20]. We implant nitrogen ions at 6 keV and a 7° angle to avoid ion channeling, resulting in a depth of 9.3 nm 3.6 nm as predicted by SRIM, at a fluence of 2 1011 cm-2. We anneal the diamond at 800 ∘C for 2 hours under a vacuum of 7.8 10-7 mbar using a custom-built furnace (Edwards T-Station 75, Tectra UHV compatible boralectric heater HTR-1001) to form NV centers [22]. To remove non-diamond carbon that may form during annealing, we clean all diamonds in a Tri-Acid solution (HNO3 65 %, HClO4 70 %, H2SO4 96 % - 1:1:1) on a hot plate set at 500 ∘C for 1 hour. We sonicate each sample sequentially in acetone (ACE) and isopropanol (IPA) before blow-drying with N2. To improve handling, we glue the diamonds onto identical silicon substrates using a thin Crystal Bond 509 layer, followed by cleaning via stirring in IPA and blow-drying with N2.
@@
Processing diamonds using Medusa 84 SiH requires an adhesion promoting layer. Following our previous work on FOx 16 [21], we deposit a 25 nm silicon layer using electron beam evaporation (Pfeiffer Classic 500 L). This step was not used for the silicon substrates. To further enhance adhesion of Medusa 84 SiH to the silicon, we eliminate moisture through hotplate drying immediately prior to resist spin coating.
@@
We apply Medusa 84 SiH (SX AR-N 8400, Allresist) using a spin coater (Süss Microtec RCD8). We mitigate edge bead formation using a tailored multi-step process. Mitigating edge bead formation here is crucial due to the small lateral dimensions of our diamonds. Medusa 84 SiH underwent a soft bake on a hot plate set at 100 ∘C for 2 minutes. In addition to the silicon adhesion promoting layer, which decreases charging during electron beam lithography [21], we apply a layer of the conductive resist ESpacer 300Z (Showa Denko) at 4000 rpm on top of Medusa 84 SiH. This additional step minimizes electrical charging, thereby enhancing the effectiveness of proximity correction during electron beam lithography (EBL). We soft bake ESpacer 300Z at 80 ∘C for 90 seconds. We note that Medusa 84 SiH is sensitive to both humidity and heat. Thus, all process steps after exposure to air were executed quickly to minimize effects on the effective writing dose. EBL was performed at 30 kV on a Raith eLiNE system (beam diameter 2 nm at 20 kV), utilizing focus-based working distance correction. We develop Medusa 84 SiH by stirring for 90 s in the developer AR 300-44 (Allresist), then rinsing with ultrapure water and cleaning with IPA for 30 s each, followed by blow-drying with N2. We note as AR 300-44 is a developer based on 2.38 % TMAH in water, the development also removes ESpacer 300Z or Electra 92. We report no residues originating from ESpacer 300Z or Electra 92 were observed during process testing. Optionally, a rinsing step in ultra pure water for 30 s - 60 s directly before development can be done to separate the removal of ESpacer 300Z / Electra 92 from the development of Medusa 84 SiH.
@@
We use three designs for EBL. First, design 1 from which we obtain the contrast curve (60 m squares and spacing). We expose each square uniformly with doses ranging from 50 C/cm2 to 5 mC/cm2, adhering to a logarithmic sequence. Design 2 features multiple 25 m squares and circles with diameters ranging from 16 m to 180 nm, spacing varying between 3 m - 6 m (as shown in Fig. 1 (a)). We perform proximity correction using Raith’s NanoPECS. This is accomplished via the dose scaling method. We employ the Monte Carlo simulation tool to estimate the proximity effect function, characterizing the radial energy distribution deposited by electrons in the resist. Most of the exposure dose results from low-energy secondary electrons generated by the high-energy electron beam during forward and backward scattering events within the resist. Utilizing the proximity function, the algorithm divides the initial design into smaller rectangles and basic polygons, such as triangles, allowing for adjustment with individual dose factors between 1 and 16 to achieve an approximately uniform dose distribution as intended by the original design. Design 3 consists of a 4 3 array of fields with 220 nanopillars (2.5 m spacing) in each field and additional markers (letters, numbers, as shown in Fig. 1 (b) and (c)). Within the array, we adjust the nanopillar diameter (180 nm, 200 nm, 220 nm, and 240 nm) and an extra dose factor (0.9, 1.0, and 1.1). We thus enable the examination of NV center spin characteristics in different pillar geometries and partly counterbalance variations in the resist layer thickness due to edge beads. We subject design 3 to proximity correction with identical parameters, however, nanopillars were not segmented and written as circles with uniform doses. As a result from proximity correction, nanopillars typically received dose factors around 14 - 16 depending on their surroundings. For our studies, we use a dose array ranging from 160 C/cm2 to 430 C/cm2 in 30 C/cm2 increments. We here present the results achieved with a base dose of 310 C/cm2, therefore the nanopillars were manufactured typically with doses around the three nominal values: 4.464 mC/cm2, 4.960 mC/cm2, or 5.456 mC/cm2.
@@
To pattern the diamond, we employ two dedicated etching processes: Initially, we remove the silicon interlayer in-between the patterned mask using a biased SF6 plasma. Subsequently, an oxygen-based ”Pillar Etch” process transfers the mask pattern into diamond [6]. For both steps, we employ an ICP-RIE (Sentech PTSA-ICP Plasma Etcher SI 500, refer to Table 2 for parameters).
@@
After structuring, we remove residual Medusa 84 SiH using buffered oxide etching (BOE) and residual silicon using potassium hydroxide (KOH). We employ Tri-Acid cleaning, stiring in IPA and blow-drying with N2 before characterizing the NV centers.
@@ 3 Process Characterization
This study reveals that Medusa 84 SiH offers performance comparable to the cutting-edge resist FOx. FOx is renowned for its effectiveness on diamond and similar substrates that require dense oxygen plasmas with significant physical etching in ICP-RIE, while allowing for nanoscale structuring. We have verified that charge dissipating resists like ESpacer 300Z or Electra 92 (Allresist) do not adversely affect Medusa 84 SiH. While an adhesion-promoting interlayer, such as silicon or titanium, is necessary for using Medusa 84 SiH, our results suggest that existing processes employing these with FOx by DuPont [6, 21, 23], and potentially other HSQ variants by EM Resist, or Applied Quantum Materials, can be readily adapted to Medusa 84 SiH. With the discontinuation of DuPont’s (formerly DowDuPont) HSQ products, our research on Medusa 84 SiH supports it as a feasible substitute, filling the gap in various research domains that heavily relied on DuPont’s FOx.
@@ 3.1 Processing Quality
Fig. 1 (a-c) presents SEM images (Hitachi SU8000) and optical microscopy images of nanopillars etched into the electronic grade diamond before mask removal (4 min pillar Etch, pillar height 430 nm). Fig. 1 (a) illustrates micrometer-scale structures that have undergone proximity correction. The sharpness of corners and edges, the uniform thickness of the resist layer (data not shown), and the lack of both overexposure and thickness steps due to dose scaling evidence effective proximity correction, which is challenging on insulating substrates such as ultrapure diamond. Fig. 1 (b) presents the nanopillar array described above, indicating good adhesion properties to the silicon interlayer and therefore a high fabrication yield. Apart from small areas in which the adhesion-layer has been scratched, the fabrication yield of the nanopillars is only negligibly impacted by adhesion failure. Fig. 1 (c) presents the investigated nanopillar field with nominal diameters of 180 nm. As a results of proximity correction, even nanopillars close to larger marker structures and at the array border are written with defined, uniform shape. Fig. 1 (d) illustrates the sidewalls of a marker structure fabricated on a test diamond for examining mask erosion. The diamond was etched for 8 minutes with the”Pillar Etch”, achieving a structure height of 915 nm 4 nm as determined by surface profilometry (Bruker DekTak XT, 1 Å vertical resolution). The fabrication involved a Medusa 84 SiH layer, originally 300 nm 10 nm thick, which was thinned down by 90 nm 20 nm with 210 nm 10 nm remaining, demonstrating its suitability for O2 plasma ICP-RIE processes (in-depth discussion in Sec. 3.3). The smooth and steep sidewall, displaying minimal roughness, indicates high writing quality, spatially homogeneous resist properties and minimal mask erosion. Fig. 1 (d) includes an inset that provides a magnified view of six nanopillars with an aspect ratio of approximately 2.4. SEM images indicate minimal roughening of the diamond surface as well as no mask redeposition effects related to Medusa 84 SiH.

### v2

@@ Abstract
We investigate the novel electron beam resist Medusa 84 SiH by Allresist GmbH (Germany) for nanostructuring of single crystal diamond and its effects on the spin properties of nitrogen vacancy (NV) centers in nanopillars as prototypes for photonic structures. We find contrast curves comparable to previously employed resists (Hydrogensilsequioxane FOx). We present a selectivity for diamond etching of 11 to 12. Using an adhesion-promoting silicon interlayer potentially enables fabrication yields of up to 96%. We measure times of up to 25 s before and after processing, demonstrating that the manufactured structures are usable for diamond-based quantum sensing.
@@
In addition, diamond is chemically inert, making it a suitable sensor material for harsh chemical environments. However, considering nanofabrication, chemical inertness poses a challenge; obtaining nanostructures typically requires plasma etching of diamond in a high-density plasma in an inductively coupled plasma reactive ion etcher (ICP-RIE). These plasmas contain oxygen and use a high bias to obtain steep sidewalls of diamond nanostructures. Such plasma conditions, however, pose high demands on the mask material used for etching; strong mask erosion might occur and consequently limit the etch depth and aspect ratio of diamond nanostructures. While metal masks fabricated in lift-off processes can withstand the plasma etching condition, they might lead to non-optimal side walls of the nanostructures. For many previous works, a mask which is obtained directly via electron beam lithography has been used: Hydrogen Silsesquioxane (HSQ)-based electron beam resists are able to withstand harsh conditions, as the chemical composition of the exposed and developed resist closely resembles SiO2 (”spin-on-glass”).
@@
A main product used in literature [4, 5] as well as in our previous work [6, 20, 21, 7] that satisfies these specifications is Flowable Oxide (FOx, DuPont, formerly DowDuPont). In 2023, DuPont notified distributors that production of FOx-related products would cease at the end of 2025. To make nanofabrication routines for diamond quantum technologies future-proof, also in the light of cost-effectiveness (upper four digit $ range per liter with minimum order quantity of one liter and shelf life of 6 months for FOx), we here present a process for a novel HSQ based resist called Medusa 84 SiH (Allresist GmbH), which has recently become commercially available. The active polymer HSQ is manufactured solely by Allresist and is not related to currently commercially available alternatives. In contrast to DuPont’s FOx, a key difference lies in the chosen solvent which is butyl acetate instead of methyl isobutyl ketone (MIBK), where the latter is considered to be possibly carcinogenic to humans, meaning Medusa 84 SiH also comes with the advantage of higher personnel safety during its application. When stored at 10 ∘C, its shelf life is analogous to DuPont’s FOx (at least 3 months from the date of purchase, extending to 6 months) and deep-freeze storage to further prolong shelf life is possible. In this work, we use this novel resist to pattern diamond nanopillars as they are typical for scanning probe sensing using color centers as a prototypical nanostructuring process. The color centers embedded in our nanostructures are individual nitrogen vacancy (NV) centers, which are the most common color centers for quantum sensing.
@@ 2 Nanofabrication Process
To make the following description of our nanofabrication process more readable, we summarize the parameters for devices, materials, and processes in tables in the supplementary material in table S.1 and S.2. In this study, we use three types of samples to evaluate the performance of Medusa 84 SiH. First, we use single crystal, single side polished silicon substrates (Microchemicals), to create accurate contrast curves and examine the writing dose-dependent etch resistivity of Medusa 84 SiH. Second, we use (100)-oriented, single crystal, chemical vapor deposited diamond (Element Six, 3.2 mm 3.2 mm (0.250 mm 0.050 mm)), for which the manufacturer specifies a moderate surface quality (Ra 30 nm) and less than 1 ppm substitutional nitrogen [N]s. We use this material to develop a consistent nanofabrication process and to study mask erosion. Third, we use a high purity (100)-oriented single crystal electronic grade diamond (Element Six), with dimensions of 4.0 mm 2.0 mm (0.047 mm 0.002 mm) ([N]s 5 ppb, [B] 1 ppb). This diamond has been polished by Delaware Diamond Knives to a surface roughness of R = 0.6 nm 0.5 nm as determined from 5 atomic force microscopy scans (5 m 5 m, Bruker FastScan Bio). With this sample, we investigate how Medusa 84 SiH influences NV centers in nanopillars, which are often used in sensing applications.
@@
We here follow published approaches to create shallow NV centers [22, 6, 20]. To obtain coherent NV centers in low-strain material, we remove polishing damage prior to nitrogen implantation by conducting a ”Stress Relief Etch” using an ICP-RIE (Oxford Instruments Plasmalab 100, for etch recipes refer to S.2) [20]. We implant nitrogen ions at 6 keV and a 7° angle to avoid ion channeling, resulting in a depth of 9.3 nm 3.6 nm as predicted by SRIM, at a fluence of 2 1011 cm-2. We anneal the diamond at 800 ∘C for 2 hours under a vacuum of 7.8 10-7 mbar using a custom-built furnace (Edwards T-Station 75, Tectra UHV compatible boralectric heater HTR-1001) to form NV centers [22]. To remove non-diamond carbon that may form during annealing, we clean all diamonds in a Tri-Acid solution (HNO3 65 %, HClO4 70 %, H2SO4 96 % - 1:1:1) on a hot plate set at 500 ∘C for 1 hour. We sonicate each sample sequentially in acetone (ACE) and isopropanol (IPA) before blow-drying with N2. To improve handling, we glue the diamonds onto identical silicon substrates using a thin Crystal Bond 509 layer, followed by cleaning via stirring in IPA and blow-drying with N2.
@@
Processing diamonds using Medusa 84 SiH requires an adhesion promoting layer. Following our previous work on FOx 16 [21], we deposit a 25 nm silicon layer using electron beam evaporation (Pfeiffer Classic 500 L). We note that thickness reduction of the layer might be feasible but has not been investigated in this work. This step was not used for the silicon substrates. To enhance adhesion of Medusa 84 SiH, we eliminate moisture through hotplate drying immediately prior to spin coating.
@@
We apply Medusa 84 SiH (SX AR-N 8400, Allresist) using a spin coater (Süss Microtec RCD8). We reduce edge bead formation using a tailored two-step process, described below. Reducing edge bead formation here is crucial due to the small lateral dimensions of our diamonds. To gradually transfer the resist from the diamond to the silicon carrier chip, we increase rotational speed slowly to 1500 rpm (acceleration 500 rpm/s, 3 s at 1500 rpm). Subsequently, we ramp to 4000 rpm (acceleration 1000 rpm/s) and maintain 4000 rpm for 30 s. We developed this two-step recipe starting from our recipe for thicker resists such as FOx 16 (higher rotational speed ( 5000 rpm) and acceleration of 2000 rpm/s in second step). Using the modified recipe lead to thicknesses ranging from 130 nm to 300 nm and patternable lateral space up to 47 % of sample area. We apply the same recipe for all 3 sample types described above. We observe variability both between and within individual diamonds, attributed to lateral inhomogeneity caused by residual edge bead formation. Medusa 84 SiH underwent a soft bake on a hot plate set at 100 ∘C for 2 minutes. In addition to the silicon adhesion promoting layer, which decreases charging during electron beam lithography [21], we apply a layer of the conductive resist ESpacer 300Z (Showa Denko) at 4000 rpm on top of Medusa 84 SiH. This additional step minimizes electrical charging, thereby enhancing the effectiveness of proximity correction during electron beam lithography (EBL). We evaluated Electra 92 (Allresist) as alternative to ESpacer 300Z. Both performed similarly well when used with Medusa 84 SiH (see table S.1 for process details). We soft bake ESpacer 300Z at 80 ∘C for 90 seconds. We note that Medusa 84 SiH is sensitive to both humidity and heat. Thus, all process steps after exposure to air were executed quickly to minimize effects on the effective writing dose. EBL was performed at 30 kV on a Raith eLiNE system (beam diameter 2 nm at 20 kV), utilizing focus-based working distance correction. We develop Medusa 84 SiH by stirring for 90 s in the developer AR 300-44 (Allresist), then rinsing with ultrapure water and cleaning with IPA for 30 s each, followed by blow-drying with N2. We note as AR 300-44 is a developer based on 2.38 % TMAH in water, the development also removes ESpacer 300Z or Electra 92. We report no residues originating from ESpacer 300Z or Electra 92 were observed during process testing. Both manufacturers suggest rinsing in ultra pure water (30 s - 60 s) to remove ESpacer 300Z / Electra 92 directly before resist development. We tested removing both during development and by rinsing prior to development without noticeable difference.
@@
We use three designs for EBL. First, design 1 from which we obtain the contrast curve (60 m squares and spacing). We expose each square uniformly with doses ranging from 50 C/cm2 to 5 mC/cm2, adhering to a logarithmic sequence. Design 2 features multiple 25 m squares and circles with diameters ranging from 16 m to 180 nm, spacing varying between 3 m - 6 m (as shown in Fig. 1 (a)). We perform proximity correction using Raith’s NanoPECS. This is accomplished via the dose scaling method. We employ the Monte Carlo simulation tool to estimate the proximity effect function, characterizing the radial energy distribution deposited by electrons in the resist. Most of the exposure dose results from low-energy secondary electrons generated by the high-energy electron beam during forward and backward scattering events within the resist. Utilizing the proximity function, the algorithm divides the initial design into smaller rectangles and basic polygons, such as triangles, allowing for adjustment with individual dose factors between 1 and 16 to achieve an approximately uniform dose distribution as intended by the original design. Design 3 consists of a 4 3 array of fields with 220 nanopillars (2.5 m spacing) in each field and additional markers (letters, numbers, as shown in Fig. 1 (b,c)). Within the array, we adjust the nanopillar diameter (180 nm, 200 nm, 220 nm, and 240 nm) and an extra dose factor (0.9, 1.0, and 1.1). We thus enable the examination of NV center spin characteristics in different pillar geometries and partly counterbalance variations in the resist layer thickness due to edge beads. We subject design 3 to proximity correction with identical parameters, however, nanopillars were not segmented and written as circles with uniform doses. As a result from proximity correction, nanopillars typically received dose factors around 14 - 16 depending on their surroundings. For our studies, we use a dose array ranging from 160 C/cm2 to 430 C/cm2 in 30 C/cm2 increments. We here present the results achieved with a base dose of 310 C/cm2, therefore the nanopillars were manufactured typically with doses around the three nominal values: 4.464 mC/cm2, 4.960 mC/cm2, or 5.456 mC/cm2.
@@
To pattern the diamond, we employ two dedicated etching processes: Initially, we remove the silicon interlayer in-between the patterned mask using a biased SF6 plasma. Subsequently, an oxygen-based ”Pillar Etch” process transfers the mask pattern into diamond [6]. For both steps, we employ an ICP-RIE (Sentech PTSA-ICP Plasma Etcher SI 500, refer to table S.2 for parameters).
@@
After structuring, we remove residual Medusa 84 SiH using buffered oxide etching (BOE) and residual silicon using potassium hydroxide (KOH). To finalize nanofabrication, we perform Tri-Acid cleaning to remove residual contamination and establish oxygen-containing surface termination followed by stirring in IPA and blow-drying with N2 before characterizing the NV centers.
@@ 3 Process Characterization
This study reveals that Medusa 84 SiH offers performance comparable to the cutting-edge resist FOx, which is renowned for its effectiveness on diamond and similar substrates that require dense oxygen plasmas with significant physical etching in ICP-RIE, while allowing for nanoscale structuring. We have verified that charge dissipating resists like ESpacer 300Z or Electra 92 (Allresist) do not adversely affect Medusa 84 SiH. While an adhesion-promoting interlayer, such as silicon or titanium, is necessary for using Medusa 84 SiH, our results suggest that existing processes employing these with FOx by DuPont [6, 21, 23], and potentially other HSQ variants by EM Resist, or Applied Quantum Materials, can be readily adapted to Medusa 84 SiH. With the discontinuation of DuPont’s HSQ products, our research on Medusa 84 SiH supports it as a feasible substitute, filling the gap in various research domains that heavily relied on DuPont’s FOx.
@@ 3.1 Processing Quality
Fig. 1 (a-c) presents SEM images (Hitachi SU8000) and optical microscopy images of nanopillars etched into the electronic grade diamond before mask removal (4 min ”Pillar Etch”, pillar height 430 nm). Fig. 1 (a) illustrates micrometer-scale structures that have undergone proximity correction. The sharpness of corners and edges, the uniform thickness of the resist layer (data not shown) and the lack of both overexposure and thickness steps due to dose scaling evidence effective proximity correction, which is challenging on insulating substrates such as ultrapure diamond. Fig. 1 (b) presents the nanopillar array described above, indicating good adhesion properties to the silicon interlayer and therefore a high fabrication yield. Apart from small areas in which the adhesion-layer has been scratched, the fabrication yield of the nanopillars is only negligibly impacted by adhesion failure. Fig. 1 (c) presents the investigated nanopillar field with nominal diameters of 180 nm. As a result of proximity correction, even nanopillars close to larger marker structures and at the array border are written with defined, uniform shape. Fig. 1 (d) illustrates the sidewalls of a marker structure fabricated on a test diamond for examining mask erosion. The diamond was etched for 8 minutes with the ”Pillar Etch”, achieving a structure height of 915 nm 4 nm as determined by surface profilometry (Bruker DekTak XT, 1 Å vertical resolution). The fabrication involved a Medusa 84 SiH layer, originally 300 nm 10 nm thick, which was thinned down by 90 nm 20 nm with 210 nm 10 nm remaining, demonstrating its suitability for O2 plasma ICP-RIE processes (in-depth discussion in Sec. 3.3). The smooth and steep sidewall, displaying minimal roughness, indicates high writing quality, spatially homogeneous resist properties and minimal mask erosion. Fig. 1 (d) includes an inset providing a magnified view of six nanopillars with an aspect ratio of and a top diameter of 214 nm 5 nm indicating slight overexposure as lithography targeted 180 nm. SEM images indicate minimal roughening of the diamond surface as well as no mask redeposition effects related to Medusa 84 SiH. Utilizing an AFM (Park XE-70), we determine the surface roughness of the etched surface (from 12 scans, 1.25 1.25 m). We find on average R = 2 nm 1 nm slightly higher than R = 0.6 nm 0.5 nm. We identify individual etch pits as primary cause increasing Rq. As such individual etch pitch are not detrimental for the process demonstrated here, we excluded scans with prominent etch pits and determine R = 0.8 nm 0.2 nm, close to R. The findings clearly indicate that RIE-ICP etching utilizing Medusa 84 SiH maintains surface quality for well-polished substrates.